A rainwater collection tank is nearly full at a level of 120 in. The water will be used at a rate that lowers the level 6 in. per day. Which equation can be used to determine the time it will take (in days) to lower the level to 3 in.?

Answer by bucky(2189)

From the problem you know that the water level needs to drop from 120 inches deep to 3 inches
deep. Therefore, the total amount of this drop is 120 - 3 inches.
.
Each day the water drops 6 inches. So if you multiply 6 times the number of days (D) you will
get the total drop. Set these two total drops equal. In equation form it becomes:
.
6*D = 120 - 3
.
Subtract the two numbers on the right side. When you do, you get that the amount of drop
in the tank is 117 inches. Substitute this amount into the equation, and the equation
becomes:
.
6*D = 117 inches
.
But we want to solve for D, the number of days that it takes for the level in the tank
to be down to 3 inches. Do this by dividing both sides of the equation by 6, the multiplier
of D. When you do, the equation is reduced to:
.
D = 117/6 = 19 + 3/6 = 19 + 1/2 = 19.5 days
.
In 19 and a half days the water will drop from 120 inches deep to 3 inches deep if the
water level drops 6 inches a day.
.
Hope this helps you to understand the problem and how to set up and solve the equation.